OpenStudy (amtran_bus):

Gas law:

OpenStudy (amtran_bus):

A mixture of 1.30 g sample of KClO3 and 1.10 g of MnO2 is decomposed by heat. 420 mL of O2 is collected. The resulting test tube and contents lost 0.50 g of mass during the RXN. The gas was collected over water Barometric pressure was 750mm, and the temp was 27C. Vapor pressure of H2) at 27 degrees C is 27mm.

OpenStudy (amtran_bus):

I know v=NRT/P

OpenStudy (amtran_bus):

@abb0t @chmvijay @aaronq

OpenStudy (amtran_bus):

Any idea?

OpenStudy (amtran_bus):

Help

OpenStudy (aaronq):

what are you trying to find?

OpenStudy (amtran_bus):

V on V=NRT/P

OpenStudy (aaronq):

why is \(P_{H_2}\) important?

OpenStudy (amtran_bus):

I have absolutely no idea. This is a lab question for this week. Sorry.

OpenStudy (aaronq):

so were assuming that the 0.50 g lost is equivalent to the amount of \(O_2\) evolved?

OpenStudy (amtran_bus):

Correct.

OpenStudy (aaronq):

okay, so \(V=\dfrac{\dfrac{0.50g}{18g/mol}*R*(27+273.15)K}{750\;mm\;Hg}\)

OpenStudy (aaronq):

just use the right R

OpenStudy (amtran_bus):

Thanks so much. Can you help me with R?

OpenStudy (aaronq):

no problem. http://en.wikipedia.org/wiki/Gas_constant

OpenStudy (amtran_bus):

Your awesome.

OpenStudy (amtran_bus):

Thanks!

OpenStudy (amtran_bus):

Then it asks to find percent error if the theoretical molar volume is 22.41. I got 0.09242 for above. *eek*

OpenStudy (amtran_bus):

@Hailey553

OpenStudy (aaronq):

damn, i noticed i told you to divide by 18 g/mol, it should've been 32 g/mol, i was thinking of water. you should fix that